# Parameters Optimization in Trajectory Planning Using Diffrentiable Convex Programing

### v1

@@ 4.2.1 Entry Trajectory Planning for Hypersonic Gliding Vehicle
$\bar{q}=k_{q}\rho^{0.5}V^{3.15}\leqslant q_{\max}$
$\bar{n}=\frac{\sqrt{L^{2}+D^{2}}}{mg_{0}}\leqslant n_{\max}$
where $k_{q}$ is the heat rate constant, and $g_{0}$ is the gravitational acceleration. The terminal velocity must satisfy $V\geq V_{\mathrm{f,min}}$. The angle and angular-rate constraints are
@@
$J_{1,\mathrm{HGV}}$ $=T_{\mathrm{ref}}\sum_{n=1}^{N}\left(2C_{1}\gamma_{\mathrm{ref}}[n]\delta\gamma[n]+2C_{2}\dot{\alpha}_{\mathrm{ref}}[n]\delta\dot{\alpha}[n]+2C_{2}\dot{\sigma}_{\mathrm{ref}}[n]\delta\dot{\sigma}[n]\right)$ (58)
$+\delta T\sum_{n=1}^{N}\left(C_{1}\gamma^{2}_{\mathrm{ref}}[n]+C_{2}\dot{\alpha}^{2}_{\mathrm{ref}}[n]+C_{2}\dot{\sigma}^{2}_{\mathrm{ref}}[n]\right)$
@@
$=J_{1,\mathrm{HGV}}+\sum_{n=1}^{N}\left(\omega_{\gamma}\delta\gamma^{2}[n]+\omega_{u}\left(\delta\dot{\alpha}^{2}[n]+\delta\dot{\sigma}^{2}[n]\right)\right)$
@@ 4.2.2 Optimization of Trust-Region Penalty Coefficients
$C_{L21}$ 1.477 $\lambda_{0}$ $0.0^{\circ}$ $\lambda_{\mathrm{f}}$ $120.0^{\circ}$ $\alpha_{\mathrm{max}}$ $40.0^{\circ}$
@@
Simulation results for different trust-region penalties are shown in Figs. 7 and 8. Only the region $\omega_{\gamma}\in[10^{1},\,10^{2}],\omega_{u}\in[10^{-1},\,10^{2}]$ is displayed, because smaller $\omega_{\gamma}$ values fail to provide sufficient regularization for the flight-path-angle state, causing SCP to diverge within the maximum iteration limit (20 iterations). Accordingly, the displayed region represents the primary domain in which SCP approaches or achieves convergence.
@@
Figure 8 presents the corresponding distribution of the cost function. The valley of minimum cost closely aligns with the region of rapid convergence, confirming that the proposed convergence metric provide a meaningful characterization of SCP performance. The red trajectory illustrates the parameter optimization update direction of the algorithm from a representative initial point. It exhibits a continuous descending trend along the cost contours, indicating that the cost function with respect to the penalty parameters possesses a smooth and differentiable structure. This observation further validates that the proposed DSCP framework is amenable to effective optimization in the parameter space. Figure 9 depicts the gradient field obtained by DSCP backpropagation, quantified by $\left\|\nabla_{\omega_{\gamma},\,\omega_{u}}\mathcal{L}_{\mathrm{HGV}}\right\|$. The gradient field exhibits a clear outward divergence from the minimum-cost region, reflecting the structure of the underlying cost landscape and confirming that DSCP accurately captures the sensitivity of the objective with respect to the trust-region penalty parameters. In contrast, for SCP configurations that fail to converge, the gradients follow the expected outward trend only in a global sense while exhibiting localized irregularities in certain regions. These deviations indicate that effective parameter optimization under DSCP still benefits from appropriately chosen initial penalty values.
@@
(a) Performance inde over DSCP iterations.
@@
(a) Case1: $\omega_{\gamma}=1.429,\omega_{u}=1.536$.
(b) Case2: $\omega_{\gamma}=1.333,\omega_{u}=0.388$.
(c) Case3: $\omega_{\gamma}=1.344,\omega_{u}=-0.199$.
(d) Case4: $\omega_{\gamma}=1.346,\omega_{u}=-0.635$.
@@ 4.3 Geometric Parameters Optimization of HGVs
The surface-to-mass ratio is a key design parameter for HGVs. Since the aerodynamic coefficients can be further tailored through airframe shaping, the present study neglects the influence of the surface-to-mass ratio on the aerodynamic coefficients themselves. Based on the nominal value $C_{S/m,0}$ reported in [31], the surface-to-mass ratio is parameterized as
@@
where $\omega_{S/m}$ denotes the scaling factor. The aim of this experiment is to identify the optimal value of $\omega_{S/m}$ within the interval $[0.75,1.25]$, thereby providing insight into the geometric design characteristics of the vehicles. To further assess the effectiveness of the optimized trust-region coefficients identified in Sec. 4.2, four representative cases of trust-region parameters are employed for the optimization of $\omega_{S/m}$. All other simulation settings are kept identical to those used in Sec. 4.2.
@@
Figures 1315 illustrate the overall optimization behavior. Figure 13(a) shows the performance index as a function of $\omega_{S/m}$ for the number of discretization intervals $N_{\mathrm{HGV}}=100$. Although the curves corresponding to different trust-region parameters do not coincide exactly, they exhibit highly similar shapes and yield nearly identical optimal values of $\omega_{S/m}$. Notably, Notably, several nonsmooth points are observed along the curves, at which the derivative of the cost function is discontinuous. As shown in Fig. 13(b), these nonsmooth locations coincide precisely with changes in the number of SCP iterations required for convergence, indicating that the observed nonsmoothness originates from the numerical behavior of the SCP procedure. Since SCP approximates a continuous-time optimal-control problem through both temporal discretization and iterative convexification, abrupt variations in the number of required iterations may occur. This nonsmooth behavior is progressively mitigated as the convergence tolerance is tightened and the number of discretization nodes is increased. This trend is confirmed by the results in Fig. 14 for $N_{\mathrm{HGV}}=200$, where the cost function becomes significantly smootherand the iteration numbers exhibit improved stability.
The statistical results in Table 3 indicate that the optimal surface-to-mass ratios for both discretization intervals are around $\omega_{S/m}^{\star}=0.93$. This suggests that the nominal airframe design is close to optimal with respect to the chosen performance index, and deviations from optimality may stem from additional geometric or aerodynamic factors that are not explicitly modeled in this study. Moreover, the trust-region coefficients optimized in Sec. 4.2 continue to exhibit favorable numerical performance, with Case 4 achieving the smallest iteration numbers across all experiments.
@@ 5 Conclusions
This paper develops the DSCP framework for learning and optimizing trajectory planning parameters in aerospace applications. By leveraging differentiable convex optimization, the proposed approach derives sensitivity relations between problem parameters and the optimal solution of SOCP based on first-order optimality conditions. These sensitivities are then combined with the chain rule and gradient backpropagation, enabling end-to-end differentiability from the optimization objective to the underlying parameters. As a result, differentiable mappings are established that support the pretraining of decision variables, the tuning of hyperparameters in trajectory optimization and SCP formulations, and parameter learning within trajectory planning pipelines. The proposed framework is evaluated in two representative scenarios: powered descent guidance and HGV entry. Optimization tasks include the terminal time optimization in powered descent, the trust-region penalty parameters tuning in SCP subproblems, and the surface-to-mass ratio optimization for HGVs. Simulation results demonstrate that the DSCP framework can reliably perform parameter optimization and prelearning of optimization variables. These findings indicate that the proposed approach provides a powerful and versatile tool for vehicle design, mission design, and hyperparameter selection in trajectory planning algorithms across a broad range of aerospace applications.
@@
5.1 details in PDG
5.2 details in HGV

### v2

@@ 4.2.1 Entry Trajectory Planning for Hypersonic Gliding Vehicle
$\bar{q}=k_{q}\rho^{0.5}V^{3.15}\leqslant\bar{q}_{\max}$
$\bar{n}=\frac{\sqrt{L^{2}+D^{2}}}{mg_{0}}\leqslant\bar{n}_{\max}$
where $k_{q}$ is the heat rate constant, and $g_{0}$ is the gravitational acceleration. The angle and angular-rate constraints are
@@
$J_{1,\mathrm{HGV}}$ $=T_{\mathrm{ref}}\sum_{n=0}^{N}\left(2C_{1}\gamma_{\mathrm{ref}}[n]\delta\gamma[n]+2C_{2}\dot{\alpha}_{\mathrm{ref}}[n]\delta\dot{\alpha}[n]+2C_{2}\dot{\sigma}_{\mathrm{ref}}[n]\delta\dot{\sigma}[n]\right)$ (58)
$+\delta T\sum_{n=0}^{N}\left(C_{1}\gamma^{2}_{\mathrm{ref}}[n]+C_{2}\dot{\alpha}^{2}_{\mathrm{ref}}[n]+C_{2}\dot{\sigma}^{2}_{\mathrm{ref}}[n]\right)$
@@
$=J_{1,\mathrm{HGV}}+\sum_{n=0}^{N}\left(\omega_{\gamma}\delta\gamma^{2}[n]+\omega_{u}\left(\delta\dot{\alpha}^{2}[n]+\delta\dot{\sigma}^{2}[n]\right)\right)$
@@ 4.2.2 Optimization of Trust-Region Penalty Coefficients
$C_{L21}$ 1.477 $\chi_{0}$ $0.0^{\circ}$ $\chi_{\mathrm{f}}$ $120.0^{\circ}$ $\alpha_{\mathrm{max}}$ $40.0^{\circ}$
@@
Simulation results for different trust-region penalties are shown in Figs. 7 and 8. Only the region $\omega_{\gamma}\in[10^{1},\,10^{2}],\omega_{u}\in[10^{-1},\,10^{2}]$ is displayed, because smaller $\omega_{\gamma}$ values fail to provide sufficient regularization for the flight-path angle, causing SCP to diverge within the maximum iteration limit (20 iterations). Accordingly, the displayed region represents the primary domain in which SCP approaches or achieves convergence.
@@
Figure 8 presents the corresponding distribution of the cost function. The valley of minimum cost closely aligns with the region of rapid convergence, confirming that the proposed convergence metric provide a meaningful characterization of SCP performance. The red trajectory illustrates the parameter optimization update direction of the algorithm from a representative initial point. It exhibits a continuous descending trend along the cost contours, indicating that the cost function with respect to the penalty parameters possesses a smooth and differentiable structure. This observation further validates that the proposed DSCP framework is amenable to effective optimization in the parameter space. Figure 9 depicts the gradient field obtained by DSCP backpropagation, quantified by $\left\|\nabla_{\omega_{\gamma},\,\omega_{u}}\mathcal{L}_{\mathrm{HGV}}\right\|$. The gradient field exhibits a clear outward divergence from the minimum-cost region, reflecting the structure of the underlying cost landscape and confirming that DSCP accurately captures the sensitivity of the objective with respect to the trust-region penalty parameters. In contrast, for SCP configurations that fail to converge, the gradients follow the expected outward trend only in a global sense while exhibiting localized irregularities in certain regions. These deviations indicate that effective parameter optimization under DSCP still benefits from appropriately chosen initial parameter values.
@@
(a) Performance index over DSCP iterations.
@@
(a) Case1: $\log_{10}(\omega_{\gamma})=1.429,\log_{10}(\omega_{u})=1.536$.
(b) Case2: $\log_{10}(\omega_{\gamma})=1.333,\log_{10}(\omega_{u})=0.388$.
(c) Case3: $\log_{10}(\omega_{\gamma})=1.344,\log_{10}(\omega_{u})=-0.199$.
(d) Case4: $\log_{10}(\omega_{\gamma})=1.346,\log_{10}(\omega_{u})=-0.635$.
@@ 4.3 Geometric Parameters Optimization of HGVs
The surface-to-mass ratio is a key design parameter for HGVs. Since the aerodynamic coefficients can be further tailored through airframe shaping, the present study neglects the influence of the surface-to-mass ratio on the aerodynamic coefficients themselves. Based on the nominal value $C_{S/m,0}$ reported in [zhangReentryTrajectoryOptimization2025], the surface-to-mass ratio is parameterized as
@@
where $\omega_{S/m}$ denotes the scaling factor. The aim of this experiment is to identify the optimal value of $\omega_{S/m}$ within the interval $[0.75,1.25]$, thereby providing insight into the geometric design characteristics of the vehicles. To further assess the effectiveness of the optimized trust-region coefficients identified, four representative cases of trust-region parameters, which are the same as Sec. 4.2, are employed for the optimization of $\omega_{S/m}$. All other simulation settings are kept identical to those used in Sec. 4.2.
@@
Figures 1315 illustrate the overall optimization behavior. Figure 13(a) shows the performance index as a function of $\omega_{S/m}$ for $N_{\mathrm{HGV}}=100$. Although the curves corresponding to different trust-region parameters do not coincide exactly, they exhibit highly similar shapes and yield nearly identical optimal values of $\omega_{S/m}$. Notably, several nonsmooth points are observed along the curves, at which the derivative of the cost function is discontinuous. As shown in Fig. 13(b), these nonsmooth locations coincide precisely with changes in the number of SCP iterations required for convergence, indicating that the observed nonsmoothness originates from the numerical behavior of the SCP procedure. Since SCP approximates a continuous-time optimal-control problem through both temporal discretization and iterative convexification, abrupt variations in the number of required iterations may occur. This nonsmooth behavior is progressively mitigated as the convergence tolerance is tightened and the number of discretization nodes is increased. This trend is confirmed by the results in Fig. 14 for $N_{\mathrm{HGV}}=200$, where the cost function becomes significantly smootherand the iteration numbers exhibit improved stability.
The statistical results in Table 3 indicate that the optimal surface-to-mass ratios for both discretization intervals are around $\omega_{S/m}^{\star}=0.92$. This suggests that the nominal airframe design is close to optimal with respect to the chosen performance index, and deviations from optimality may stem from additional geometric or aerodynamic factors that are not explicitly modeled in this study. Moreover, the trust-region coefficients optimized in Sec. 4.2 continue to exhibit favorable numerical performance, with Case 4 achieving the smallest iteration numbers across all experiments.
@@ 5 Conclusions
This paper develops the differentiable sequential convex programming (DSCP) framework for learning and optimizing trajectory planning parameters in aerospace applications. By leveraging differentiable convex optimization, the proposed approach derives sensitivity relations between problem parameters and the optimal solution of SOCP based on first-order optimality conditions. These sensitivities are then combined with the chain rule and gradient backpropagation, enabling end-to-end differentiability from the optimization objective to the underlying parameters. As a result, differentiable mappings are established that support the pretraining of decision variables, the tuning of hyperparameters in trajectory optimization and SCP formulations, and parameter learning within trajectory planning pipelines. The proposed framework is evaluated in two representative scenarios: powered descent guidance and HGV entry. Optimization tasks include the terminal time optimization in powered descent, the trust-region penalty parameters tuning in SCP subproblems, and the surface-to-mass ratio optimization for HGVs. Simulation results demonstrate that the DSCP framework can reliably perform parameter optimization and prelearning of optimization variables. These findings indicate that the proposed approach provides a powerful and versatile tool for vehicle design, mission design, and hyperparameter selection in trajectory planning algorithms across a broad range of aerospace applications.
@@
5.1 Details in Powered Descent Guidance
In this subsection, the specific settings and constructions of subproblems for PDG are provided. For convenience, the subscripts "PDG" are omitted in the following equations. The subproblem for PDG is expressed as
$\min$ $-m[N]+\omega_{\Gamma}\sum_{n=0}^{N-1}\left\|\Gamma[n]\right\|_{2}^{2}$ (65)
$\mathrm{s.t.}$ $\mathrm{boundary\,constraints}:$
$\boldsymbol{x}[0]=\boldsymbol{x}_{0},\boldsymbol{r}[N]=\boldsymbol{r}_{\mathrm{f}},\boldsymbol{v}[N]=\boldsymbol{v}_{\mathrm{f}}$
$\mathrm{discretized\,dynamics\,constraints}\,(n\in[0,N-1]):$
$\left(\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}+\boldsymbol{I}\right)\delta\boldsymbol{x}[n]+\left(\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}-\boldsymbol{I}\right)\delta\boldsymbol{x}[n+1]+\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}\delta\boldsymbol{u}[n]+\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}\delta\boldsymbol{u}[n+1]$
$=(\boldsymbol{x}_{\text{ref}}[n+1]-\boldsymbol{x}_{\text{ref}}[n])-T_{\text{ref}}\boldsymbol{f}_{\text{ref}}[n]+\boldsymbol{\Gamma}[n]$
$\mathrm{thrust\,constraints}\,(n\in[0,N]):$
$u_{\min}\leq\left\|\boldsymbol{u}_{\mathrm{ref}}[n]\right\|_{2}+\frac{\boldsymbol{u}_{\mathrm{ref}}^{\top}[n]}{\left\|\boldsymbol{u}_{\mathrm{ref}}[n]\right\|_{2}}\left(\boldsymbol{u}[n]-\boldsymbol{u}_{\mathrm{ref}}[n]\right)\leq u_{\max}$
$\sqrt{u_{y}^{2}[n]+u_{z}^{2}[n]}\leq\tan(\eta_{\max})u_{x}[n]$
$\mathrm{path\,constraints}\,(n\in[0,N]):$
$\sqrt{r_{y}^{2}[n]+r_{z}^{2}[n]}\leq\tan(\beta_{\max})r_{x}[n]$
where $\boldsymbol{\Gamma}$ is the virtual control vector and $\omega_{\Gamma}=10000$ is the weighting coefficient for the virtual control. The dynamics Jacobians are given by
$\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}=\begin{bmatrix}\boldsymbol{0}_{3\times 3}&\boldsymbol{I}_{3\times 3}&\boldsymbol{0}_{3\times 1}\\ \boldsymbol{0}_{3\times 3}&\boldsymbol{0}_{3\times 3}&-\frac{\boldsymbol{u}}{m}\\ \boldsymbol{0}_{1\times 3}&\boldsymbol{0}_{1\times 3}&0\end{bmatrix},\quad\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}=\begin{bmatrix}\boldsymbol{0}_{3\times 3}\\ \frac{\boldsymbol{I}_{3\times 3}}{m}\\ -\frac{\boldsymbol{u}^{\top}}{I_{\mathrm{sp}}g_{0}\left\|\boldsymbol{u}\right\|_{2}}\end{bmatrix}$ (66)
The initial guess of the SCP procedure for PDG is provided by a simplified problem given by
$\min$ $\sum_{n=0}^{N}\left\|\boldsymbol{u}[n]\right\|_{2}^{2}$ (67)
$\mathrm{s.t.}$ $\boldsymbol{x}[0]=\boldsymbol{x}_{0},\boldsymbol{r}[N]=\boldsymbol{r}_{\mathrm{f}},\boldsymbol{v}[N]=\boldsymbol{v}_{\mathrm{f}}$
$\boldsymbol{r}[n+1]-\boldsymbol{r}[n]=\frac{T}{2}\left(\boldsymbol{v}[n+1]+\boldsymbol{v}[n]\right)$
$\boldsymbol{v}[n+1]-\boldsymbol{v}[n]=\frac{T}{2}\left(\frac{\boldsymbol{u}[n+1]}{m[n+1]}+\frac{\boldsymbol{u}[n]}{m[n]}+2\boldsymbol{g}_{E}\right),m[n]=m_{0}$
In the above problem, the variation of mass and the thrust constraints are disregarded, so the resulting trajectory does not satisfy the constraints of PDG. In the early stages of the iteration, the virtual control can effectively compensate for the deficiencies of the initial trajectory.
5.2 Details in Entry Trajectory Optimization
In this subsection, we detail the formulation and setup of the subproblems associated with entry trajectory planning for HGVs. For simplicity, the subscript "HGV" is omitted in the subsequent equations. The subproblem can be written as
[…]
$\mathrm{s.t.}$ $\mathrm{boundary\,constraints}:$
$\boldsymbol{x}[0]=\boldsymbol{x}_{0},\boldsymbol{x}[N]=\boldsymbol{x}_{\mathrm{f}}(\mathrm{except\,for},V[N]),V[n]\geq V_{\mathrm{f,\min}}$
$\mathrm{discretized\,dynamics\,constraints}\,(n\in[0,N-1]):$
$\left(\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}+\boldsymbol{I}\right)\delta\boldsymbol{x}[n]+\left(\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{x}}\right|_{\text{ref}}-\boldsymbol{I}\right)\delta\boldsymbol{x}[n+1]+\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}\delta\boldsymbol{u}[n]+\frac{T_{\text{ref}}}{2}\left.\frac{\partial\boldsymbol{f}}{\partial\boldsymbol{u}}\right|_{\text{ref}}\delta\boldsymbol{u}[n+1]$
$=(\boldsymbol{x}_{\text{ref}}[n+1]-\boldsymbol{x}_{\text{ref}}[n])-T_{\text{ref}}\boldsymbol{f}_{\text{ref}}[n]$
$\mathrm{path\,constraints}\,(n\in[0,N]):$
$\bar{p}[n]+\left.\frac{\partial\bar{p}}{\partial r}\right|_{\text{ref}}\delta r[n]+\left.\frac{\partial\bar{p}}{\partial V}\right|_{\text{ref}}\delta V[n]\leq\bar{p}_{\max}$
$\bar{q}[n]+\left.\frac{\partial\bar{q}}{\partial r}\right|_{\text{ref}}\delta r[n]+\left.\frac{\partial\bar{q}}{\partial V}\right|_{\text{ref}}\delta V[n]\leq\bar{q}_{\max}$
$\bar{n}[n]+\left.\frac{\partial\bar{n}}{\partial r}\right|_{\text{ref}}\delta r[n]+\left.\frac{\partial\bar{n}}{\partial V}\right|_{\text{ref}}\delta V[n]+\left.\frac{\partial\bar{n}}{\partial\alpha}\right|_{\text{ref}}\delta\alpha[n]\leq\bar{n}_{\max}$
$\mathrm{control\,angle\,constraints}\,(n\in[0,N]):$
$\alpha_{\min}\leq\alpha[n]\leq\alpha_{\max},\,\sigma_{\min}\leq\sigma[n]\leq\sigma_{\max}$
$\dot{\alpha}_{\min}\leq\dot{\alpha}[n]\leq\dot{\alpha}_{\max},\,\dot{\sigma}_{\min}\leq\dot{\sigma}[n]\leq\dot{\sigma}_{\max}$
The dynamics Jacobians and the state-dependent differentials of the path constraints are omitted here for brevity and can be found in [wang2017constrained]. The initial trajectory for all entry simulations is calculated using a fixed control profile, which is given by
$u_{1}(t)=\frac{\frac{\pi}{9}-\alpha_{0}}{t_{\mathrm{f0}}}$ (69)
$u_{2}(t)=-\,\frac{2\pi}{3}\,\frac{\sigma_{\max}}{t_{\mathrm{f0}}}\cos\!\left(\frac{2\pi t}{t_{\mathrm{f0}}}\right)-\frac{1}{6}\,\frac{\sigma_{\max}}{t_{\mathrm{f0}}}.$
where $t_{\mathrm{f0}}=2500s$ is the final time for the initial reference trajectory, $\alpha_{0}$ is the initial angle of attack, and $\sigma_{\max}$ is the bank angles, respectively. These settings guarantee that the initial reference controls satisfy the angle and angular-rate constraints in Eq.(56).